Geoff the Giraffe walked 117 miles, beginning on Sunday morning and finishing on Monday evening of the following week. Each day, he walked 1 mile farther than the day before. How many miles did Geoff walk each day?
Answer by ankor@dixie-net.com(22740)

Geoff the Giraffe walked 117 miles, beginning on Sunday morning and finishing on Monday evening of the following week. Each day, he walked 1 mile farther than the day before. How many miles did Geoff walk each day?
:
let d = no. mi walked the first day, Sunday
Assuming he walked 9 days, which would be next Monday
Day Mi walked
s - d
m - d+1
t - d+2
w - d+3
t - d+4
f - d+5
s - d+6
s - d+7
m - d+8
--------
tot 9d + 36 = 117
9d = 117 - 36
d = 81/9
d = 9 mi walked on the first Sunday, You can find the no. of mi the rest of the days